OpenStudy (anonymous):

Let f(x) = 2x + 2. Solve f-1(x) when x = 4.

OpenStudy (anonymous):

@_lilb_to_smooth_4_U_

OpenStudy (anonymous):

replace 4 with x so it looks like f(x)=2(4)+2

OpenStudy (anonymous):

so the answer is 10? can u help me re word 5 more problems like this?:)

OpenStudy (anonymous):

sure and are there answer choices or are you doing all these like that

OpenStudy (anonymous):

1) Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). 2) Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). 3) Let f(x) = 12 over the quantity of 4 x + 2. Find f(-1). 4) Let f(x) = x2 - 3x - 7. Find f(-3). 5) Let f(x) = 7x - 13. Find f-1(x).

OpenStudy (anonymous):

@ftflyingtiger can u just re word them like u did in the above problem please:)

OpenStudy (anonymous):

@meghana333666999

OpenStudy (anonymous):

@emmigrace222

OpenStudy (anonymous):

1) Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). There ask for f (x^2-7x-9)(-1)

OpenStudy (anonymous):

all you are doing is plugging in so if they asked g(f(1) you would multiply g by what f is

OpenStudy (anonymous):

im still confused:(

OpenStudy (zarkon):

if this is your question Let \(f(x) = 2x + 2\). Solve \(f^{-1}(x)\) when \(x = 4\). then the answer is 1 since \(f(1)=2\cdot 1+2=4\)